Sol Summary – February 13th

Crew 221 Sol Summary Report 13Feb-2020

Sol:11

Summary Title: Real Mars Experience

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Today the crew woke up and had pancakes prepared by the crew commander and with the olive oil we got from the last supply mission from Earth. Finally, they do not stick to the pan anymore and were delicious!

In the morning we were under the gun to get the robot ready as there were hookups with the software. All bugs were solved by 13:00, leaving one hour before the EVA.

The girls EVA team, containing the crew commander M.Grulich, crew journalist H. Kalucha and crew GreenHab Officer J. Todd who mainly worked on the robot, got ready for the EVA. They decided to test the robot just in front of the hab door to make sure it was working. It did but the wheels came loose, and it had to be fixed in the field. This was a real challenge, requesting several tools from the team support inside the hab.

Once all issues were fixed the EVA team headed for the rover but realized that Spirit was not plugged in. After radio communication, it was decided to leave one team member at the hab and go with permission with two crew members on Curiosity. Crew commander and crew GreenHab officer drove to Steambed Connector and walked to Candor Chasma and tested the robot successfully in a more challenging environment.

Once back the team had a debrief discussing the communication and confusion of information that lead to one person not joining the EVA mission and realized that any piece of information from mission support should and shall in the future be reviewed and read in more detail, as we cannot rely on mission support on Mars.

The evening was spent with a lovely salad prepared by crew HSO Rawan Alshammari and will be concluded with a games night.

Anomalies in work: None

Update mouse in hab: no new mice caught

Weather: sunny light winds

Crew Physical Status: Crew some sore throat, but seems to improve

Reports to be filed:

Sol Summary
Photo Report
Operations Report
Journalist Report
GreenHab Report
EVA Request
EVA Report 13

Support Requested: None

Greenhab Report – February 13th

Crew 221 GreenHab Report 13FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. The door opened today from 10:14 pm to 2:22 pm.

Average temperatures: 18°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 16.51 gallons

Daily water usage for research and/or other purposes: 0.52 gallon

Water in Blue Tank: 152.3 gallons

Time(s) of watering for crops: 2

Changes to crops: I repotted the avocado into a larger pot with fresh soil and perlite and removed the prickly weed. Didn’t water the narcissus, so the pots can dry out.

Narrative: Today was a garden maintenance day. I repotted the avocado (which is doing surprisingly well!) into a much larger pot and removed the horrid prickly weed that was growing around it. I added some perlite to a few of the newly potted plants and tidied up the radish and carrot beds. Fresh carrots for dinner, yum!

More news on the Martian corn, it’s now growing out of all but 1 Martian pot (apparently Martian clay isn’t ideal for corn growing), even a little resilient sprout out of clay mixed with earth soil.

Our mungbeans are trying to sprout, but I don’t think the Martian climate is agreeing with them. We also harvested some of the older greens, in preparation for working on those plant beds tomorrow.

Harvest: 26g Chard, 7g Rocket, 8g Spinach, 42g Carrots, 12g Lettuce.

Support/supplies needed: NA

Other Notes: Can I clarify; the tap I can use is the big antique looking one at the tunnel entrance to the GreenHab with the big orange tape that says “do not use”? Also, Shannon, you would like me to pick all the plants in the two-tier grey pot by the door (i.e. the chard, spinach and rocket on the top, and the sprouting spinach on the bottom) or just the top pot?

Operations Report – February 13th

Crew 221 Operations Report 13Feb2020

SOL: 11

Name of person filing report: Luis D. Monge

Non-nominal systems:
Notes on non-nominal systems:

Generator: on (at 8:20 pm)
Hours run: 12:40
From what time last night: 7:40 pm
To what time this morning: 7:00 am
List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before the generator is run at night) 73%

Diesel Reading – 75%
Station Propane Reading – 48%
Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons
Water Meter: 148 units
Water (static tank) – 180 gals
Static to Loft Pump used – no
Water in GreenHab – 152.3 gals (water is lower than estimated, please refer to Greenhab report)
Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: no
Hours: (before EVA) : 130.4
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging)
Currently charging: yes

Opportunity rover used: no
Hours: n/a
Beginning charge: n/a
Ending charge: n/a
Currently charging: yes

Curiosity rover used: yes
Hours: 137.2
Beginning charge: 100%
Ending charge: 78%
Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used
Reason for use: not used
Oil Added? not used
ATV Fuel Used: Gals not used
# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment set to finish tomorrow, now watered with 2l/day. watering of the plants carried out as per the procedure

Summary of ScienceDome operations: one oven is operating in the lab, the incubator is non-functional, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly completed, radio communication experiment completed.

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report.

Questions, concerns, and requests to Mission Support: Spirit was disconnected; EVA crew reconnected it.

EVA Report – February 13th

Crew 221 EVA Report 13FEB-2020

EVA #13

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was to test our Martian rover, MOUSIE. EVA #7 selected sites of interest in Candor Chasma, that were inaccessible to astronauts and featured some interesting rock stratification that we wanted to sample. Today’s EVA aimed to test our rover in these sites.

Start time: 2:30 pm
End time: 4:32 pm

Narrative: Today is the day to finally test our robot! Our crack team of roboticists (Hermani – Crew Journalist, Jessica – GreenHab Officer, Maria – Commander) has spent the better part of the last two weeks building MOUSIE (the Mars Overland and Underground Scientific Investigative Explorer). After a morning filled with last-minute debugging, involving four separate computers and five different crewmates, MOUSIE was finally ready for deployment.

We were headed to Candor Chasma, to send MOUSIE where no Mars astronaut had gone (or could go) before. Candor Chasma features some interesting rock overhangs that the astronauts can’t access in their bulky EVA suits, but are a perfect site to test MOUSIE. We did some initial tests out the front of the Hab, in close proximity to our brains’ trust inside if something went wrong.

A last-minute wheel reattachment (quite the task in bulky EVA gloves) and we were ready to go. Upon discovering one of the intended rovers was not charged, Hermani offered to remain behind, and Maria and Jessica set off for Candor Chasma in the other rover.

Tests in the Chasma were a success! MOUSIE was able to access rock samples under the overhangs using her little robotic arm, though she did struggle on the more sandy patches at the bottom of the Chasma. MOUSIE 2.0 might need some more intense overland tires!

Destination: Walking along Stream Bed Connector and Candor Chasma.

Coordinates (use UTM NAD27 CONUS): 0520285/4251221

Participants: Jessica Todd (GreenHab Officer), Maria Grulich (Commander), Hemani Kalucha (Journalist)

Road(s) and routes per MDRS Map: Drove along Entrance Road then north along Cow Dung Road (0110). Parked at entrance to Stream Bed Connector and walked along with it to Candor Chasma.

Mode of travel: Driving and Walking

Operations Report – February 12th

MDRS Supplemental Operations Report 12-FEB-2020 (Updated)

Name of person filing report: Shannon Rupert

Reason for Report: Routine Update/Return to campus

Non-nominal systems: Generator

Action is taken for non-nominal systems: Currently on a rental generator.

We determined that the generating unit has failed in the generator.
Because it has been replaced before (and it about half the price of a
new generator) and the engine has over 18K hours, we will be replacing the generator. The plan is to keep the rental for another month while finding a 10K replacement generator that is extended use and is compatible with the power system’s autostart. The power system itself is nominal.

Generator check, note if oil and coolant added: Checked levels on the
rental, all are nominal at this time.

ScienceDome AC unit, on and operational: Still not operational.

Solar— running nominally, solar throughout the day and generator
backup usually on 12 hours a night

Diesel Reading – not checked
Propane Reading – not checked
Ethanol Free Gasoline – not checked

Water (director’s tank) – not checked
Water (static tank) – not checked
Water (GreenHab) – not checked
Water (loft) – not checked
Static to Loft Pump used – n/a
Water Meter: n/a

Toilet tank emptied: n/a

Deimos rover used: Yes. Repairs completed. This rover has new
batteries which need to be used 30 times before they are fully
charged. This rover can now be used by crews as long as:

1. Rover is not taken out as the only rover
2. Rovers still return to station when the first one reaches 60 %
3. Keep an eye on the charge throughout the EVA. If you see a quick drop in the charge (say from 70 to 40 % in a few seconds) come home immediately.
4. You can take the rover out with just one person in each rover
while this rover is fully charging its batteries.

Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Sojourner rover used: nominal
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: nominal
Hours:
Beginning charge:
Ending charge:

Currently charging: yes. This rover now has the metal parts that
ripped the spacesuits covered. It is a good plan and we will cover
the others as soon as possible.

Opportunity rover used: Still has the brake issue and should not be
used until I work things out with the repair shop who failed to make
the repair. Other than the brake issue, it is nominal.

Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Curiosity rover used: nominal.
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Notes and action took on rovers: I changed the power cord that was
being used to charge Opportunity. It was missing its grounding prong.
In addition, I secured both charging ports at the RAM so the plug is
protected.

Monthly check and fill of rover batteries: Not done at this time.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.1 is the only ATV
operating nominally. The 300 is still in town. 350.2 has a safety
issue with the damaged handlebars and currently does not have a
battery. 350.2 and the Honda have low batteries

Reason for use: Started and checked for operations
Oil Added? No
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used: n/a

Notes and action took on ATVs: I have the jumper cables from the Hab and will work on them tomorrow. I’m also going to try to adjust the handlebars on the damaged rover.

HabCar running? If used, why, where? Yes, used for trips to town. It
has the wrong battery in it (it’s the battery is currently on the
generator) and it is filthy. Needs new jumper cables and its plastic
shovel for mud

CrewCar running, If used why, where? Currently in GJ. Needs to be
returned to complete repairs that were not completed last time it was
in the shop (Atila, please confirm)

Campus-wide inspection, if an action is taken, what and why? Mostly nominal.
Small diesel spill at the tank has gotten larger and will need
remediation. The leak has been stopped. Rockpile in front of Hab has
been removed. Looks great. Someone has created a new and incorrect route to deliver water than goes between the Hab and the observatory and solar array. In the future, the correct route needs to be used.

Evidence of rodents, where and action is taken: Two rodents caught in Hab and released. Crews need to make sure all food is in cupboards and the kitchen is clean every night. There is evidence that a rodent was in my trailer, but apparently they have moved out.

Summary of internet: All three accounts nominal. Suggest replacement of wifi box in Hab (I can never remember which is the router and which is the modem)

Summary of suits and radios: Four yellow radios were added to the
crew supply. Two were from the ScienceDome originally so only two new ones were added. Radios are the most replaced item on campus. Did not check the suits.

Summary of Hab operations: Nothing to report at this time as the crew is in sim

Summary of Outpost operations: The discharge plumbing to the intern
trailer has been installed. However, the replacement of dirt is still a
work in progress and needs to be completed.

Summary of GreenHab operations: New seed bank can and additional seeds
were added to the GreenHab. Nothing else to report as the crew is in
sim at this time.

Summary of ScienceDome operations: Dual split still not completed.
The incubator is non-functional due to a failure in its digital chip.
Crews will need to use the oven for an incubator. Additional soil
needs to be banked up and over the cement foundation outside to
protect the interior floor from water damage.

Summary of RAM operations: The wiring issue at the charging stations has been completed. There is new damage to the top of the outer airlock door and no one seems to know how it happened.

Summary of any observatory issues: A cord was replaced by Atila and
Peter and both observatories are in good shape.

Summary of health and safety issues: The stove was installed with part of the changeover from natural gas to propane not completely done.

The crew identified the problem and it has been corrected. Stove
operations are now nominal.

List and summary of projects: Now that I am back on campus, I will be
working with Marc and Mike from Crew 216 to rework the MDRS map.
I’m excited about this project!

Questions, concerns, supplies needed and requests: Nothing at this time.

Journalist Report – February 13th

Crew 221 Journalist Report 13Feb2020

Author: Hemani Kalucha, Crew Journalist

Sol 11

Title: Almost at the End

This morning, we got the program of the rover to function! In the afternoon, the all-girl (Maria, Hemani, Jess) EVA team headed out to test the rover at Candor Chasma. We got some adorable pictures of the rover picking up samples under overhang rocks. After celebrating our last EVA, the crew plans to spend one of our last evenings together with games and good conversation.

EVA Report – February 12th

Crew 221 EVA Report 12Feb 2020
EVA # 11

Author: Maria Grulich (Commander)

Purpose of EVA: We will be performing the deployment of a dipole antenna and document the whole process. The antenna will be deployed near to the habitat with radio instruction from the team that performed the procedure in the previous day.

Start time: 10:40 am
End time: 12:30 pm

Narrative:

Crew commander M. Grulich, Crew Journalist H. Kalucha, and Crew Greenhab Officer J. Todd went out to see how easy it is for an untrained crew to install the antenna of Crew Engineer L. Monge. The weather was great, and the team found a metal pool for the first installation site, which was still in communication range to the hab where L. Monge was crew support.

The installation went smoothly within a record time of 15 min so the team decided to go for another site where there was no support structure. The location was a rock formation where the pole could be locked to install the antenna, showing that such a concept could be done on Mars.

On the way back, the team of three carried all the equipment back and lost “Fuzz” our mascot that we took out for outreach videos. This resulted in a rescue walk back as Fuzz feels like a real crew member now and was actually found. Everyone was relieved.

The team is preparing an MDRS dance video and the EVA team spent 15 min on shooting some parts of it so stay tuned. The EVA crew was back at 12:30, meeting some Martians fixing the diesel leak.

Destination: close to the habitat

Coordinates (use UTM NAD27 CONUS): close to the habitat

Participants: Maria Grulich (Commander), Hemani Kalucha (Crew Journalist), Jessica Todd (GreenHab Officer)

Road(s) and routes per MDRS Map: None

Mode of travel: Walking

Vehicles used: N/A